# Filtering with Confidence: When Data Augmentation Meets Conformal Prediction

### v1

@@ Abstract
With promising empirical performance across a wide range of applications, synthetic data augmentation appears a viable solution to data scarcity and the demands of increasingly data-intensive models. Its effectiveness lies in expanding the training set in a way that reduces estimator variance while introducing only minimal bias. Controlling this bias is therefore critical: effective data augmentation should generate diverse samples from the same underlying distribution as the training set, with minimal shifts. In this paper, we propose conformal data augmentation, a principled data filtering framework that leverages the power of conformal prediction to produce diverse synthetic data while filtering out poor-quality generations with provable risk control. Our method is simple to implement, requires no access to internal model logits, nor large-scale model retraining. We demonstrate the effectiveness of our approach across multiple tasks, including topic prediction, sentiment analysis, image classification, and fraud detection, showing consistent performance improvements of up to 40% in $F_{1}$ score over unaugmented baselines, and 4% over other filtered augmentation baselines.
@@ 1 Introduction
Synthetic data augmentation refers to a set of machine learning techniques and heuristics designed to artificially expand a training dataset Shorten & Khoshgoftaar (2019); Taqi et al. (2018). As noted by Huang et al. (2022), practitioners have long relied on augmenting inputs with perturbed versions of the original data—both to enhance model robustness to small perturbations and based on the general intuition that “more data is always better.” With the emergence of advanced foundation models capable of generating remarkably high-quality synthetic data however (from images (Ho et al., 2020; Karras et al., 2017, 2019; Ramesh et al., 2022; Rombach et al., 2022), to text (Brown et al., 2020; Li et al., 2022; Touvron et al., 2023), or molecular structures (Jin et al., 2018; Shi et al., 2020)), synthetic data generation has experienced renewed interest. Such approaches promise significant practical advantages, particularly in reducing the time, cost, and effort involved in augmenting datasets through additional data collection and annotation Nadas et al. (2025). Synthetic data augmentation has already demonstrated promising empirical results across a wide range of applications. In natural language processing, it has been effectively used for model fine-tuning on small datasets and in low-resource language settings (Feng et al., 2020; Yang et al., 2019; Li et al., 2020; Mahamud et al., 2023; Wang et al., 2022), as well as for knowledge base construction (Li et al., 2024b) , etc. In computer vision, it has shown benefits in tasks such as image classification (He et al., 2016) and object detection (Bochkovskiy et al., 2020), etc.
@@
From a theoretical perspective, much is still to determine about the benefits of synthetic data. Recent theoretical insights from Huang et al. (2022); Nakada et al. (2024) have begun characterizing the effect of synthetic oversampling in certain regimes on estimator error bounds. Intuitively, synthetic oversampling should work well if it manages to enlarge the training set, reducing estimator variance whilst only incurring a slightly increased bias. Synthetic augmentation methods thus face a fundamental tension. On one hand, generated samples should closely follow the distribution of the original data to minimize bias—typically requiring using lower variability in the generation (or a “low temperature") to ensure that the generated data remains faithful to the original. On the other hand, synthetic samples need to be sufficiently diverse and decorrelated to be treated effectively as independent observations, a goal typically achieved by increasing generation variability (e.g., raising the temperature parameter) Havrilla et al. (2024).
@@
Despite the current enthusiasm for synthetic data sampled from generative AI models, no principled approach has yet been proposed to determine this trade-off systematically Jordon et al. (2022). In fact, current methods for generating synthetic data exhibit limited flexibility in their handling of samples with varying levels of quality. To adapt the loss to various levels of synthetic data quality, some techniques, such as the approach by Jaine et al. Jain et al. (2024) or that of Nakada et al. Nakada et al. (2024), introduce hyperparameters to control the weights placed on the reconstruction errors corresponding to the original data and the synthetic data respectively, effectively putting less emphasis on the synthetic data if its quality is too low. But these methods are inherently inflexible and treat all generated datapoints similarly. In particular, these methods are unable to distinguish between good and bad synthetic examples, thereby effectively discarding all synthetic data points from distributions that produce mixtures of high-quality and low-quality outputs Alaa et al. (2022); Rajeswar et al. (2023); Ravuri & Vinyals (2019) . Finer methods, capable of operating effectively in high-variability ("high-temperature”) regimes and explicitly distinguishing high-quality generated samples from poor ones, are still lacking.
@@
    Provable guarantees of control of our procedure over the number of poor quality samples accepted in the augmented data using approximate conditional coverage in our setting (Section 3.2). Our method adapts the framework of Gibbs et al. (2025); Cherian et al. (2024) to provide robust, condition-specific quality guarantees.
@@ 2 Background: Synthetic Data Generation and Filtering
Figure 1: Illustration of the workflow in clinical disease prediction. Data augmentation candidate outputs from the generative model $h$ (GPT-4.1 nano in this example) are filtered by a quality predictor trained on $\mathcal{D}_{train}$ with a threshold calibrated by $\mathcal{D}_{\text{calib}}$. The retained output preserves the meaning of “common cold,” while the discarded output does not correspond to the intended symptom.
@@
Formally, let $X_{i}\in\mathcal{D}$ denote an observed data point, $Y_{i}\in\mathcal{Y}$ denote additional sample meta information (such as labels captions) which we might want to condition upon in our generating procedure. We assume that the data point $X_{i}$ is sampled from a true underlying distribution $h^{\star}$ that depends on the context/label: $X_{i}\sim h^{\star}(C_{i},Y_{i})$ where $C_{i}$ represents the latent context. Intuitively, $Y_{i}$ encodes observable attributes such as class labels or side information, and $C_{i}$ captures hidden structure or nuisance variation specific to the dataset at hand, and that is not directly observed but decides how $X_{i}$ is realized. To synthesize new instances from the same distribution, we generate $K$ alternative versions of $X_{i}$ by reusing $X_{i}$ as proxy for the latent context $C_{i}$:
@@
where $h$ denotes the generative model conditioned explicitly on the observed data point and features ($X_{i}$, $Y_{i}$), and $\tau$ is a temperature parameter controlling the model’s generation variability. Thus, the generative model $h$ serves as an approximation to the true distribution $h^{\star}$, replacing the inaccessible latent context $C_{i}$ with observable surrogates $X_{i},Y_{i}$.
@@
Although effective, synthetic data from generative models can be noisy or distributionally shifted Feng et al. (2021); Kumar et al. (2020) — particularly when increasing the temperature $\tau$—, potentially reducing downstream performance. Various approaches, such as prompt engineering, direct generative modeling, retrieval-based methods, and filtering strategies (e.g., human evaluation, similarity metrics, classification-based filtering), have been proposed to improve synthetic data quality Alaa et al. (2022); Lewis et al. (2020); Liu et al. (2023). However, these filtering methods critically depend on accurate and oftentimes expensive quality metrics (such as human evaluation), which remain challenging Ding et al. (2024); Rajeswar et al. (2023). With increased generation diversity, ensuring quality becomes critical.
We propose revisiting here a simple filtration technique, as proposed in Islam et al. (2024); Kang et al. (2021); Li et al. (2024a). These methods all operate on the following premise: low-quality generations should be filtered out. Let $\mathcal{A}:\Omega\times\Omega\times\mathcal{Y}\to\mathbb{R}$ be a measure of a generated sample’s quality. Ideally, $\mathcal{A}$ should quantify the degree of deviation of the generation from the underlying data distribution. Filtering-based methods choose to remove generated examples for which $\mathcal{A}(G_{ik},X_{i},Y_{i})<\lambda$, for a user-defined threshold $\lambda$. The threshold $\lambda$ should be neither too low (to avoid content of low quality), nor too high (to avoid generating trivial rephrasings).
@@ 3.1 Controlling the number of wrong inclusions
We consider the gold standard quality scores ${\mathbf{A}}_{i}=({A}_{ik})_{k=1}^{K}$ and the corresponding surrogate scores ${\mathbf{\hat{A}}}_{i}=(\hat{A}_{ik})_{k=1}^{K}$ for the generations in the calibration data. We define the filtered set at surrogate level $s$ by the notation: $\mathcal{S}(\mathbf{\hat{A}}_{i},s)=\{G_{ik}:\hat{A}_{ik}\geq s\}$. Let $\mathcal{L}_{\lambda}(\mathcal{S}(\mathbf{\hat{A}}_{i},s),\mathbf{A}_{i})$ denote a loss function that measures the quality of filtered output compared to the ground truth $\mathbf{A}_{i}$. For instance, we may define $\mathcal{L}_{\lambda}(\mathcal{S}(\mathbf{\hat{A}}_{i},s),\mathbf{A}_{i})$ to be the number of generations $\hat{A}_{ik}$ with surrogate score at least $s$ but whose gold-standard scores $A_{ik}$ are below the nominal quality threshold $\lambda$:
@@
where $\rho$ is a hyperparameter that represents the tolerance on the loss, or the maximal number of “false discoveries” per sample that we are willing to allow. In other words, we define the non-conformity score $S(\hat{\mathbf{A}}_{i},\mathbf{A}_{i})$ as the minimal threshold $s$ such that the filtered set $\mathcal{S}(\hat{\mathbf{A}}_{i},s)$ contains only all the generations for $X_{i}$ with surrogate score $\hat{\mathbf{A}}_{ik}>s$, and at most $\rho$ of these generations have gold-standard scores $\mathbf{A}_{ik}<\lambda$.
In this paper, we formulate the problem of filtering generations based on imperfect surrogate quality scores $\hat{\mathbf{A}}$ as a calibration problem: we need to select the surrogate filtering threshold $s$ in a data-driven manner so as to ensure that $\mathbb{P}(\mathcal{L}_{\lambda}(\mathcal{S}(\mathbf{\hat{A}}_{i_{0}};s_{i_{0}}),\mathbf{A}_{i_{0}})\leq\rho)\geq 1-\alpha$ for all $i_{0}\in{\mathcal{D}_{\text{aug}}}$ with some user-specified confidence level $\alpha\in(0,1)$. To this end, we propose leveraging conformal prediction (CP) Vovk et al. (2005); Angelopoulos et al. (2022) for risk control. Conformal methods provide finite-sample, distribution-free guarantees by calibrating predictions using a hold-out validation set (see Appendix B for a more in-depth review). In our setting, we use the distribution of the scores $S(\mathbf{\hat{A}}_{i},\mathbf{A}_{i})$ to correctly calibrate our rejection threshold to ensure retaining quality content. Letting $\hat{s}_{i_{0}}$ be the output of the conformal prediction algorithm for each $\mathbf{X}_{i_{0}}$ (see the explicit formula in equation 7 in Appendix), we will solely accept generated examples with $\hat{A}_{i_{0}k}>\hat{s}_{i_{0}}$.
@@ 3.2 Conditional Conformal Risk Control
While this setup is intuitive, one could argue that, like $\lambda$, the surrogate threshold $s$ might just as well be chosen using data splitting – making the conformal prediction step appear unnecessary. However, our setting is more challenging: the difficulty of the filtering problem varies across samples, and fixed validation-based thresholds cannot adapt to this heterogeneity. To address this, we incorporate sample-specific information and apply conditional conformal prediction, allowing the filtering procedure to adapt to the hardness of each instance and thereby provide more reliable control.
While conformal prediction can act as a wrapper around any method, it is a well-established fact that it is impossible to get conditional results Barber et al. (2021). To address this, we adopt the relaxation proposed by Gibbs et al. (2025), which designs a prediction set that satisfies the guarantee over a specified function class $\mathcal{F}$:
$\mathbb{E}\left[f(X_{i_{0}})\cdot\left(\mathbf{1}\{\mathcal{L}_{\lambda}(\mathcal{S}(\mathbf{\hat{A}}_{i_{0}};s_{i_{0}}),\mathbf{A}_{i_{0}})\leq\rho)\}-(1-\alpha)\right)\right]=0,\text{ for all }f\in\mathcal{F}.$ (3)
@@
where $\mathcal{F}_{W}$ is the RKHS function class associated with $W$. The intercept $\beta$ guarantees the marginal coverage, while the RKHS term $f_{W}$ enables flexible, smooth calibration of the conformity scores $\{S_{i}\}_{i\in\mathcal{D}_{calib}}$ against covariates. As the following lemma shows, the resulting cutoff $\hat{s}_{i_{0}}$ for each data $i_{0}\in\mathcal{D}_{aug}$ satisfies the conditional guarantee over a localization region around $X_{i_{0}}$.
Lemma 1
Consider the function class $\mathcal{F}$ as defined in Equation 4, and assume $\mathcal{D}_{\text{calib}}\bigcup\mathcal{D}_{\text{aug}}$ are i.i.d. . Suppose $\mathcal{L}_{\lambda}(\cdot,\cdot)$ is monotone (i.e. for any sets $\mathcal{S}_{i_{0}}^{1}\subseteq\mathcal{S}_{i_{0}}^{2}$, it must be the case that $\mathcal{L}_{\lambda}(\mathcal{S}_{i_{0}}^{1},\mathbf{A}_{i_{0}})\leq\mathcal{L}_{\lambda}(\mathcal{S}_{i_{0}}^{2},\mathbf{A}_{i_{0}})$) and $\mathcal{L}_{\lambda}(\emptyset,\cdot)=0$. Assume $W(x,\cdot)$ defines a density with respect to each $x\in\Omega$, and sample $X_{i_{0}}^{\prime}\mid X_{i_{0}}=x\sim W(x,\cdot)$. Then for all $f\in\mathcal{F}$, $i_{0}\in\mathcal{D}_{\text{aug}}$,
$\mathbb{P}\left(\mathcal{L}_{\lambda}(\mathcal{S}(\mathbf{\hat{A}}_{i_{0}};\hat{s}_{i_{0}}),\mathbf{A}_{i_{0}})\leq\rho\mid X^{\prime}_{i_{0}}=x_{i_{0}}^{\prime}\right)=1-\alpha-\frac{\gamma\mathbb{E}[\hat{f}^{\hat{s}_{i_{0}}}_{W}(x_{i_{0}}^{\prime})]}{\mathbb{E}[W(X_{i_{0}},x_{i_{0}}^{\prime})]},$
@@
Due to the infinite dimensionality of the RKHS class, the achieved coverage departs from the nominal level $1-\alpha$ by a gap of $\frac{-\gamma\mathbb{E}[\hat{f}^{\hat{s}_{i_{0}}}_{W}(x_{i_{0}}^{\prime})]}{\mathbb{E}[W(X_{i_{0}},x_{i_{0}}^{\prime})]}$. from the nominal level $1-\alpha$. However, this coverage gap is estimable, and can be quantified using the procedure proposed in Gibbs et al. (2025). The proof of Lemma 1 is shown in Appendix C.2
In practice, we rely on the fast implementation of this approximate conditional CP algorithm as provided in Anonymous (2026), which provides a fast alternative to the original algorithm of Gibbs et al. (2025).
@@
1:Reference evaluation AA; surrogate evaluation $\hat{A}$; calibration dataset $\mathcal{D}_{\text{calib}}=\{(X_{i},(G_{ik})_{k=1}^{K},Y_{i})\}$ augmentation dataset $\mathcal{D}_{\text{aug}}=\{(X_{i},(G_{ik})_{k=1}^{K}),Y_{i})\}$; quality level $\lambda$; loss function $\mathcal{L}$; contamination allowance $\rho$
2:Compute the reference score: ${A}_{ik}=\mathcal{A}(G_{ik},X_{i},Y_{i}),\ \forall i\in\mathcal{D}_{\text{calib}}$.
3:Compute the surrogate score: ${\hat{A}}_{ik}\leftarrow\hat{\mathcal{A}}((G_{ik})_{k=1}^{K},X_{i},Y_{i}),\ \forall i\in\mathcal{D}_{\text{calib}}\cup\mathcal{D}_{\text{aug}}$.
4:Compute the non-conformity score associated with loss $\mathcal{L}$ and $\rho$, denoted as $S(\mathbf{\hat{A}}_{i},\mathbf{A}_{i})$, according to Equation 2, $\forall i\in\mathcal{D}_{\text{calib}}$.
5:for $i\in D_{\text{aug}}$ do
6:  Fit conditional conformal prediction to find $\hat{s}_{i}$
7:  Select generations : $\mathcal{S}(\mathbf{\hat{A}}_{i},\hat{s}_{i})=\{G_{ik}:\hat{A}_{ik}\geq\hat{s}_{i}\}$
8:end for
9:
10:The selected generations $\{\mathcal{S}_{i}:i\in\mathcal{D}_{\text{aug}}\}$.

### v2

@@ Abstract
With promising empirical performance across a wide range of applications, synthetic data augmentation appears a viable solution to data scarcity and the demands of increasingly data-intensive models. Its effectiveness lies in expanding the training set in a way that reduces estimator variance while introducing only minimal bias. Controlling this bias is therefore critical: effective data augmentation should generate diverse samples from the same underlying distribution as the training set, with minimal shifts. In this paper, we propose conformal data augmentation, a principled data filtering framework that leverages the power of conformal prediction to produce diverse synthetic data while filtering out poor-quality generations with provable risk control. Our method is simple to implement, requires no access to internal model logits, nor large-scale model retraining. We demonstrate the effectiveness of our approach across multiple tasks, including topic prediction, sentiment analysis, image classification, and fraud detection, showing consistent performance improvements of up to 40 percentage points (pp) in $F_{1}$ score over unaugmented baselines, and 4 pp over other filtered augmentation baselines.
@@ 1 Introduction
Synthetic data augmentation refers to a set of machine learning techniques and heuristics designed to artificially expand a training dataset (Taqi et al., 2018; Shorten and Khoshgoftaar, 2019). As noted by Huang et al. (2022), practitioners have long relied on augmenting inputs with perturbed versions of the original data—both to enhance model robustness to small perturbations and based on the general intuition that “more data is always better.” With the emergence of advanced foundation models capable of generating high-quality synthetic data however (from images (Karras et al., 2017, 2019; Ho et al., 2020; Ramesh et al., 2022; Rombach et al., 2022), to text (Brown et al., 2020; Li et al., 2022; Touvron et al., 2023), or molecular structures (Jin et al., 2018; Shi et al., 2020)), synthetic data generation has experienced renewed interest. Such approaches promise significant practical advantages, particularly in reducing the time, cost, and effort involved in augmenting datasets through additional data collection and annotation (naduaș2025synthetic). Synthetic data augmentation has already demonstrated promising empirical results across a wide range of applications. In natural language processing, it has been effectively used for model fine-tuning on small datasets and in low-resource language settings (Yang et al., 2019; Feng et al., 2020; Li et al., 2020; Wang et al., 2022; Mahamud et al., 2023), as well as for knowledge base construction (Li et al., 2024b). In computer vision, it has shown benefits in tasks such as image classification (He et al., 2016; Li et al., 2025) and object detection (Bochkovskiy et al., 2020).
@@
From a theoretical perspective, many questions remain open regarding the benefits of synthetic data. Recent theoretical insights from (Huang et al., 2022; Nakada et al., 2024) have begun characterizing the effect of synthetic oversampling in certain regimes on estimator error bounds. Intuitively, synthetic oversampling should work well if it manages to enlarge the training set, reducing estimator variance whilst only incurring a slightly increased bias. Synthetic augmentation methods thus face a fundamental tension. On one hand, generated samples should closely follow the distribution of the original data to minimize bias—typically requiring using lower variability in the generation (or a “low temperature") to ensure that the generated data remains faithful to the original. On the other hand, synthetic samples need to be sufficiently diverse and decorrelated to be treated effectively as independent observations, a goal typically achieved by increasing generation variability (e.g., raising the temperature parameter) (Havrilla et al., 2024).
@@
Despite the current enthusiasm for synthetic data sampled from generative AI models, no principled approach has yet been proposed to determine this trade-off systematically (Jordon et al., 2022). In fact, current methods for generating synthetic data exhibit limited flexibility in their handling of samples with varying levels of quality. To adapt the loss to various levels of synthetic data quality, some techniques, such as the approach by Jain et al. (2024) or that of Nakada et al. (2024), introduce hyperparameters to control the weights placed on the reconstruction errors corresponding to the original data and the synthetic data respectively, effectively putting less emphasis on the synthetic data if its quality is too low. But these methods are inherently inflexible and treat all generated data points similarly. In particular, these methods are unable to distinguish between good and bad synthetic examples, thereby effectively discarding all synthetic data points from distributions that produce mixtures of high- and low-quality outputs (Ravuri and Vinyals, 2019; Alaa et al., 2022). Finer methods, capable of operating effectively in high-variability (“high-temperature”) regimes and explicitly distinguishing high-quality generated samples from poor ones, are still lacking.
@@
    Provable guarantees of control of our procedure over the number of poor quality samples accepted in the augmented data using approximate conditional coverage in our setting (Section 3.2). Our method adapts the framework of Cherian et al. (2024); Gibbs et al. (2025) to provide robust, condition-specific quality guarantees.
@@ 2 Background: Synthetic Data Generation and Filtering
Figure 1: Illustration of the workflow in clinical disease prediction. Data augmentation candidate outputs from the generative model $h$ (GPT-4.1 nano in this example) are filtered by a quality predictor trained on $\mathcal{D}_{\text{train}}$ with a threshold calibrated by $\mathcal{D}_{\text{calib}}$. The retained output preserves the meaning of “common cold,” while the discarded output does not correspond to the intended symptom.
@@
Formally, let $X_{i}\in\mathcal{D}$ denote an observed data point, $Y_{i}\in\mathcal{Y}$ denote additional sample meta information (such as labels or captions) which we might want to condition upon in our generating procedure. For instance, in sentiment analysis, $X_{i}$ may be a text review and $Y_{i}$ its sentiment label. We assume that the data point $X_{i}$ is sampled from a true underlying distribution $h^{\star}$ that depends on the context/label: $X_{i}\sim h^{\star}(C_{i},Y_{i})$ where $C_{i}$ represents the latent context. Intuitively, $Y_{i}$ encodes observable attributes such as class labels or side information, and $C_{i}$ captures hidden structure or nuisance variation specific to the dataset at hand, and that is not directly observed but decides how $X_{i}$ is realized. To synthesize new instances from the same distribution, we generate $K$ alternative versions of $X_{i}$ by reusing $X_{i}$ as proxy for the latent context $C_{i}$:
@@
where $h$ denotes the generative model conditioned explicitly on the observed data point and features ($X_{i}$, $Y_{i}$), and $\tau$ is a temperature parameter controlling the model’s generation variability. Thus, the generative model $h$ serves as an approximation to the true distribution $h^{\star}$, replacing the inaccessible latent context $C_{i}$ with observable surrogates $(X_{i},Y_{i})$.
@@
Although effective, synthetic data from generative models can be noisy or distributionally shifted (Kumar et al., 2020; Feng et al., 2021) — particularly when increasing the temperature $\tau$—, potentially reducing downstream performance. Various approaches, such as prompt engineering, direct generative modeling, retrieval-based methods, and filtering strategies (e.g., human evaluation, similarity metrics, classification-based filtering), have been proposed to improve synthetic data quality (Lewis et al., 2020; Alaa et al., 2022; Liu et al., 2023). However, these filtering methods critically depend on accurate and oftentimes expensive quality metrics (such as human evaluation), which remain challenging (Ding et al., 2024). With increased generation diversity, ensuring quality becomes critical.
We propose revisiting here a simple filtration technique, as proposed in Kang et al. (2021); Islam et al. (2024); Li et al. (2024a). These methods all operate on the following premise: low-quality generations should be filtered out. Let $\mathcal{A}:\Omega\times\Omega\times\mathcal{Y}\to\mathbb{R}$ be a measure of a generated sample’s quality. Ideally, $\mathcal{A}$ should quantify the degree of deviation of the generation from the underlying data distribution. Filtering-based methods choose to remove generated examples for which $\mathcal{A}(G_{ik},X_{i},Y_{i})<\lambda$, for a user-defined threshold $\lambda$. The threshold $\lambda$ should be neither too low (to avoid content of low quality), nor too high (to avoid trivial rephrasings).
@@ 3.1 Controlling the number of wrong inclusions
We consider the gold standard quality scores ${\mathbf{A}}_{i}=({A}_{ik})_{k=1}^{K}$ and the corresponding surrogate scores ${\mathbf{\hat{A}}}_{i}=(\hat{A}_{ik})_{k=1}^{K}$ for the generations in the calibration data. We define the filtered set at surrogate level $s$ by the notation: $\mathcal{S}(\mathbf{\hat{A}}_{i},s)=\{G_{ik}:\hat{A}_{ik}\geq s\}$. Let $\mathcal{L}_{\lambda}(\mathcal{S}(\mathbf{\hat{A}}_{i},s),\mathbf{A}_{i})$ denote a loss function that measures the quality of filtered output compared to the ground truth $\mathbf{A}_{i}$. For instance, we may define $\mathcal{L}_{\lambda}(\mathcal{S}(\mathbf{\hat{A}}_{i},s),\mathbf{A}_{i})$ to be the number of generations $G_{ik}$ with surrogate score $\hat{A}_{ik}$ at least $s$ but whose gold-standard scores $A_{ik}$ are below the nominal quality threshold $\lambda$:
@@
where $\rho$ is a hyperparameter defining the loss tolerance, or the maximum allowed “false discoveries" per sample. In other words, we define the non-conformity score $S(\hat{\mathbf{A}}_{i},\mathbf{A}_{i})$ as the minimal threshold $s$ such that the filtered set $\mathcal{S}(\hat{\mathbf{A}}_{i},s)$ contains only all the generations for $X_{i}$ with surrogate score $\hat{\mathbf{A}}_{ik}>s$, and at most $\rho$ of these generations have gold-standard scores $\mathbf{A}_{ik}<\lambda$.
In this paper, we formulate the problem of filtering synthetic generations based on imperfect surrogate quality scores $\hat{\mathbf{A}}$ as a calibration problem: we need to select the surrogate filtering threshold $s$ in a data-driven manner so as to ensure that $\mathbb{P}(\mathcal{L}_{\lambda}(\mathcal{S}(\mathbf{\hat{A}}_{i_{0}};s_{i_{0}}),\mathbf{A}_{i_{0}})\leq\rho)\geq 1-\alpha$ for all $i_{0}\in{\mathcal{D}_{\text{aug}}}$ with some user-specified confidence level $\alpha\in(0,1)$. To this end, we propose leveraging conformal prediction (CP) (Vovk et al., 2005; Angelopoulos and Bates, 2021) for risk control. Conformal methods provide finite-sample, distribution-free guarantees by calibrating predictions using a hold-out validation set (see Appendix B for a more in-depth review). In our setting, we use the distribution of the scores $S(\mathbf{\hat{A}}_{i},\mathbf{A}_{i})$ to correctly calibrate our rejection threshold to ensure retaining quality content. Letting $\hat{s}_{i_{0}}$ be the output of the conformal prediction algorithm for each $\mathbf{X}_{i_{0}}$ (see the explicit formula in equation 7 in the Appendix), we will solely accept generated examples with $\hat{A}_{i_{0}k}>\hat{s}_{i_{0}}$.
@@ 3.2 Conditional Conformal Risk Control
One could argue that, like $\lambda$, the surrogate threshold $s$ might just as well be chosen using data splitting – making the conformal prediction step appear unnecessary. However, our setting is more challenging: the difficulty of the filtering problem varies across samples, and fixed validation-based thresholds cannot adapt to this heterogeneity. To address this, we incorporate sample-specific information and apply conditional conformal prediction, allowing the filtering procedure to adapt to the hardness of each instance and thereby provide more reliable control.
While conformal prediction can act as a wrapper around any method, it is a well-established fact that it is impossible to get conditional results (Foygel Barber et al., 2021). To address this, we adopt the relaxation proposed by Gibbs et al. (2025), which designs a prediction set that satisfies the guarantee over a specified function class $\mathcal{F}$:
$\begin{split}&\quad\mathbb{E}\left[f(X_{i_{0}})\left(\mathbf{1}\{\mathcal{L}_{\lambda}(\mathcal{S}(\mathbf{\hat{A}}_{i_{0}};s_{i_{0}}),\mathbf{A}_{i_{0}})\leq\rho)\}-(1-\alpha)\right)\right]\\ &=0,\quad\text{for all }f\in\mathcal{F}.\end{split}$ (3)
@@
where $\mathcal{F}_{W}$ is the RKHS function class associated with $W$. The intercept $\beta$ guarantees the marginal coverage, while the RKHS term $f_{W}$ enables flexible, smooth calibration of the conformity scores $\{S_{i}\}_{i\in\mathcal{D}_{\text{calib}}}$ against covariates. As the following lemma shows, the resulting cutoff $\hat{s}_{i_{0}}$ for each data $i_{0}\in\mathcal{D}_{\text{aug}}$ satisfies the conditional guarantee over a localization region around $X_{i_{0}}$.
Lemma 3.1.
Consider the function class $\mathcal{F}$ as defined in Equation 4, and assume $\mathcal{D}_{\text{calib}}\bigcup\mathcal{D}_{\text{aug}}$ are i.i.d. Suppose $\mathcal{L}_{\lambda}(\cdot,\cdot)$ is monotone (i.e. for any sets $\mathcal{S}_{i_{0}}^{1}\subseteq\mathcal{S}_{i_{0}}^{2}$, it must be the case that $\mathcal{L}_{\lambda}(\mathcal{S}_{i_{0}}^{1},\mathbf{A}_{i_{0}})\leq\mathcal{L}_{\lambda}(\mathcal{S}_{i_{0}}^{2},\mathbf{A}_{i_{0}})$) and $\mathcal{L}_{\lambda}(\emptyset,\cdot)=0$. Assume $W(x,\cdot)$ defines a density with respect to each $x\in\Omega$, and sample $X_{i_{0}}^{\prime}\mid X_{i_{0}}=x\sim W(x,\cdot)$. Then for all $f\in\mathcal{F}$, $i_{0}\in\mathcal{D}_{\text{aug}}$,
$\mathbb{P}\left(\mathcal{L}_{\lambda}(\mathcal{S}(\mathbf{\hat{A}}_{i_{0}};\hat{s}_{i_{0}}),\mathbf{A}_{i_{0}})\leq\rho\mid X^{\prime}_{i_{0}}=x_{i_{0}}^{\prime}\right)$
$=1-\alpha-\frac{\gamma\mathbb{E}[\hat{f}^{\hat{s}_{i_{0}}}_{W}(x_{i_{0}}^{\prime})]}{\mathbb{E}[W(X_{i_{0}},x_{i_{0}}^{\prime})]},$
@@
Due to the infinite dimensionality of the RKHS class, the achieved coverage departs from the nominal level $1-\alpha$ by a gap of $\frac{-\gamma\mathbb{E}[\hat{f}^{\hat{s}_{i_{0}}}_{W}(x_{i_{0}}^{\prime})]}{\mathbb{E}[W(X_{i_{0}},x_{i_{0}}^{\prime})]}$. from the nominal level $1-\alpha$. However, this coverage gap is estimable, and can be quantified using the procedure proposed in Gibbs et al. (2025). The proof of Lemma 3.1 is shown in Appendix C.2
In practice, we rely on the fast implementation of this approximate conditional CP algorithm as provided in (Jung et al., 2025), which provides a fast alternative to the original algorithm of Gibbs et al. (2025).
@@
0: Reference evaluation AA; surrogate evaluation $\hat{A}$; calibration dataset $\mathcal{D}_{\text{calib}}=\{(X_{i},(G_{ik})_{k=1}^{K},Y_{i})\}$; augmentation dataset $\mathcal{D}_{\text{aug}}=\{(X_{i},(G_{ik})_{k=1}^{K}),Y_{i})\}$; quality level $\lambda$; loss function $\mathcal{L}$; badness allowance $\rho$
1: Compute the reference score: $A_{ik}\leftarrow\mathcal{A}(G_{ik},X_{i},Y_{i}),\ \forall i\in\mathcal{D}_{\text{calib}}$.
2: Compute the surrogate score: ${\hat{A}}_{ik}\leftarrow\hat{\mathcal{A}}((G_{ik})_{k=1}^{K},X_{i},Y_{i}),\forall i\in\mathcal{D}_{\text{calib}}\cup\mathcal{D}_{\text{aug}}$.
3: Compute the non-conformity score associated with loss $\mathcal{L}$ and $\rho$, denoted as $S(\mathbf{\hat{A}}_{i},\mathbf{A}_{i})$, according to Equation 2, $\forall i\in\mathcal{D}_{\text{calib}}$.
4:for $i\in D_{\text{aug}}$ do
5:  Fit conditional conformal prediction to find $\hat{s}_{i}$
6:   Select generations: $\mathcal{S}(\mathbf{\hat{A}}_{i},\hat{s}_{i})=\{G_{ik}:\hat{A}_{ik}\geq\hat{s}_{i}\}$.
7:end for
7: The selected generations $\{\mathcal{S}_{i}:i\in\mathcal{D}_{\text{aug}}\}$.